Using laws of exponents, simplify and write the answer in exponential form:   (2^20 ÷ 2^15) × 2^3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of dividing the powers with same base = a^m÷ a^n= a^m - n
(2^20÷ 2^15) can be simplified as,= (2)^20 - 15 = 2^5
Then,
By the rule of multiplying the powers with same base = a^m× a^n= a^m + n
2^5× 2^3can be simplified as,
= (2)^5 + 3
= 2^8

Thanks